The sum of three times a number and 7 more than the number is the same as the difference between -11 and twice the number. What is the number?
Answer by stanbon(75887)

The sum of three times a number and 7 more than the number is the same as the difference between -11 and twice the number. What is the number?
------------
Let the number be "x".
EQUATION:
3x+(x+7) = 2x--11
3x+x+7 = 2x+11
4x+7 = 2x+11
2x=4
x = 2
-------------
Checking:
3*2 + 9 = 4+11
15 = 15
==============
Cheers,
Stan H.